A NOTE ON MAXIMUM VARIANCE OF ORDER STATISTICS FROM SYMMETRIC POPULATIONS

NICKOS PAPADATOS

Department of Mathematics, Section of Statistics and Operations Research, University of Athens, Panepistemiopolis, 157 84 Athens, Greece

(Received March 22, 1995; revised October 2, 1995)

Abstract. The maximum variance of order statistics from a symmetrical parent population is obtained in terms of the population variance. The proof is based on a suitable representation for the variance of order statistics in terms of the parent distribution function.

Key words and phrases: Order statistics, maximum variance, symmetric populations.